OpenStudy (anonymous):

if the mass A is 1 kg and the mass of B is 6.0 kg, determine the acceleration of the system and the tension of the string if the coefficient of friction between A and the table is 0.8. (pulley problem) I got 25 m/s as my answer but I'm not sure it's correct

OpenStudy (abmon98):

Lets Represent Mass A as M1 and Mass B as M2 okay 

OpenStudy (abmon98):

 Apply Newton Law to each of the two masses for m2:\[m2g-T=m2a\] for m1 \[T(tension)-μ*m1*g=m1*a\] 60-T=6a T-80=a -20=6a+a a=2.8 -T=6(2.8)-60 T=42.8N